# Dynamical stability of evolutionarily stable strategy in asymmetric games

### v2

@@ Appendix B ESS under Definition 2e for bihypermatrix games can only be pure
${\sf B}[\mathbf{p}_{1};\mathbf{\hat{y}}^{\gamma_{x}}]={\sf B}[\mathbf{p}_{2};\mathbf{\hat{y}}^{\gamma_{x}}]={\sf B}[\mathbf{\hat{x}};\mathbf{\hat{y}}^{\gamma_{x}}]~{}\text{and}~{}{\sf C}[\mathbf{q}_{1};\mathbf{\hat{x}}^{\nu_{y}}]={\sf C}[\mathbf{q}_{2};\mathbf{\hat{x}}^{\nu_{y}}]={\sf C}[\mathbf{\hat{y}};\mathbf{\hat{x}}^{\nu_{y}}].$ (87)
Note that for the two-strategy case, we have $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((\hat{x},1-\hat{x}),(\hat{y},1-\hat{y}))$. Hence the neighbourhood vector can be written as, $(\mathbf{x},\mathbf{y})\equiv(\mathbf{\hat{x}}+\bm{\delta}_{x},\mathbf{\hat{y}}+\bm{\delta}_{y})\equiv((\hat{x}+\delta_{x},1-\hat{x}-\delta_{x}),(\hat{y}+\delta_{y},1-\hat{y}-\delta_{y}))$. We want to cover the entire neighborhood around the internal fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})$; therefore, $\delta_{x}$ and $\delta_{y}$ can be varied independently within the infinitesimal range, and their signs can be anything—either positive or negative. Also, since the payoff is linear in its arguments, the following property—which can be easily proved using the expansion, ${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]=\sum_{i,j_{1},\cdots,j_{\gamma_{x}}}\hat{x}_{i}y_{j_{1}}\cdots y_{j_{\gamma_{x}}}{\sf B}[\mathbf{p}_{i};\mathbf{q}_{j_{1}}\cdots\mathbf{q}_{j_{\gamma_{x}}}]$—comes handy:
@@
To show that Definition 2e cannot be satisfied by a mixed state, we need to argue that ${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]+{\sf C}\left[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]$ cannot be positive everywhere in the entire neighbourhood. To this end, we expand the terms in Definition 2e to arrive at,
@@
Now, to the leading order, the R.H.S. of Eq. (LABEL:eqn:bihyp_ESS_s) is $-(b_{1}+c_{1})\delta_{x}\delta_{y}$ which for any given values of $b_{1}$ and $c_{1}$ cannot be positive for all infinitesimal (positive or negative) values of $\delta_{x}$ and $\delta_{y}$. For similar reason, in case $-(b_{1}+c_{1})\delta_{x}\delta_{y}=0$, the next order term in $\delta_{x}$ and $\delta_{y}$ of the R.H.S. of Eq. (LABEL:eqn:bihyp_ESS_s) leads to same conclusion, and so on. In conclusion, mixed states cannot be ESS under Definition 2e for bihypermatrix games. We remark the line of proof adopted herein can be extended to the case of mixed ESS under Definition 1e.

### v3

@@ Appendix C ESS under Definition 2e for bihypermatrix games can only be pure
${\sf B}[\mathbf{p}_{1};\mathbf{\hat{y}}^{\gamma_{x}}]={\sf B}[\mathbf{p}_{2};\mathbf{\hat{y}}^{\gamma_{x}}]={\sf B}[\mathbf{\hat{x}};\mathbf{\hat{y}}^{\gamma_{x}}]\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ {\sf C}[\mathbf{q}_{1};\mathbf{\hat{x}}^{\nu_{y}}]={\sf C}[\mathbf{q}_{2};\mathbf{\hat{x}}^{\nu_{y}}]={\sf C}[\mathbf{\hat{y}};\mathbf{\hat{x}}^{\nu_{y}}].$ (92)
Note that for the two-strategy case, we have $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((\hat{x},1-\hat{x}),(\hat{y},1-\hat{y}))$. Hence the neighbourhood vector can be written as, $(\mathbf{x},\mathbf{y})\equiv(\mathbf{\hat{x}}+\bm{\delta}_{x},\mathbf{\hat{y}}+\bm{\delta}_{y})\equiv((\hat{x}+\delta_{x},1-\hat{x}-\delta_{x}),(\hat{y}+\delta_{y},1-\hat{y}-\delta_{y}))$. We want to cover the entire neighborhood around the internal fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})$; therefore, $\delta_{x}$ and $\delta_{y}$ can be varied independently within the infinitesimal range, and their signs can be anything—either positive or negative. Also, since the payoff is linear in its arguments, the following property—which can be easily proved using the expansion, ${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]=\sum_{i,j_{1},\cdots,j_{\gamma_{x}}}\hat{x}_{i}y_{j_{1}}\cdots y_{j_{\gamma_{x}}}{\sf B}[\mathbf{p}_{i};\mathbf{q}_{j_{1}}\cdots\mathbf{q}_{j_{\gamma_{x}}}]$—comes in handy:
@@
To show that Definition 2e cannot be satisfied by a mixed state, we need to argue that ${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]+{\sf C}\left[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]$ cannot be positive everywhere in the entire neighbourhood. To this end, we expand the terms in Definition 2e to arrive at
@@
Now, to the leading order, the R.H.S. of Eq. (LABEL:eqn:bihyp_ESS_s) is $-(b_{1}+c_{1})\delta_{x}\delta_{y}$ which for any given values of $b_{1}$ and $c_{1}$ cannot be positive for all infinitesimal (positive or negative) values of $\delta_{x}$ and $\delta_{y}$. For similar reason, in case $-(b_{1}+c_{1})\delta_{x}\delta_{y}=0$, the next order term in $\delta_{x}$ and $\delta_{y}$ of the R.H.S. of Eq. (LABEL:eqn:bihyp_ESS_s) leads to same conclusion, and so on. In conclusion, mixed states cannot be ESS under Definition 2e for bihypermatrix games. We remark the line of the proof adopted herein can be extended to the case of mixed ESS under Definition 1e.